OpenStudy (anonymous):

select the inequality that corresponds to the given graph. number line with a closed circle on negative 3 and one third and 2 with shading in between |2x - 4| > 2 |4x + 2| > 10 |4x - 6| less than or greater to 14 |6x + 4| less than or greater to 16

OpenStudy (texaschic101):

I can tell you this....from your solution, a closed circle indicates that an equal sign is present. Therefore, answer choices 1 and 2 are not the correct answers. I am assuming the last two answer choices are less then or EQUAL. |4x - 6| <= 14 -14 <= 4x - 6 <= 14 -14 + 6 <= 4x <= 14 + 6 -8 <= 4x <= 20 -8/4 <= x <= 20/4 -1/2 <= x < = 5 (not this one) |6x + 4| <= 16 -16 <= 6x + 4 <= 16 -16 - 4 <= 6x <= 16 - 4 - 20 <= 6x <= 12 -20/6 <= x <= 12/6 - 3 1/3 <= x <= 2 closed circle on -3 1/3, closed circle on 2, shading in between